HIGH RESOLUTION INFRARED SPECTROSCOPY OF JET-COOLED PHENYL RADICAL IN THE GAS PHASE

\underline{ERIN N. SHARP-WILLIAMS}, MELANIE A. ROBERTS, DAVID J. NESBITT, JILA, National Institute of Standards and Technology and University of Colorado, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309.

Phenyl radical (C_{6}H_{5}) is one of the most important reactive intermediates, as it is formed from the homolytic cleavage of a CH bond in benzene (C_{6}H_{6}), and hence it plays a central role in the combustion of fossil fuels that are typically rich in aromatics. We recently recorded the first high resolution infrared spectra of jet-cooled phenyl radical in the gas phase. This was obtained by direct absorption laser spectroscopy in a slit-jet discharge supersonic expansion of a phenyl halide precursor (C_{6}H_{5}X, i.e. C_{6}H_{5}I and C_{6}H_{5}Br) diluted in a Neon/Helium gas mixture. We observed an A-type band, which arises from a fundamental excitation of the out-of-phase symmetric CH stretch (ν_{19}). The unambiguous assignment of the rotational structure in this band to C_{6}H_{5} is facilitated by comparing 2-line combination differences with the Fourier transform microwave (FTM) and direct absorption millimeter-wave (mm-wave) measurements of the ground state by McMahon et al.^{a}. A least-squares fit to an asymmetric top Hamiltonian of the rotationally-resolved vibrational band is done to determine upper-state rotational constants and a gas-phase band origin (ν_{0}) of 3071.8904 (10) cm^{−1}. This is in very good agreement with the value of 3071 cm^{−1} for the out-of-phase symmetric CH stretch of phenyl reported by Friderichsen et al.^{b} from matrix isolation studies, which indicates a surprisingly small red shift due to the low-temperature argon environment.

aR. J. McMahon, M. C. McCarthy, C. A. Gottlieb, J. B. Dudek, J. F. Stanton and P. Thaddeus, Ap. J. 590, L61 (2003).

bA. V. Friderichsen, J. G. Radziszewski, M. R. Nimlos, P. R. Winter, D. C. Dayton, D. E. David and G. B. Ellison, J. Am. Chem. Soc. 123, 1977 (2001).